Solve.
(s+1/4)(s+1/5)

Answer by ewatrrr(10682)

 
Hi
In order to 'solve' an Equation, it needs to be an equation.
If You meant:
(s+1/4)(s+1/5)= 0
(s+1/4)=0 then s = -1/4
(s+1/5)= 0 s = -1/5
(s+1/4)(s+1/5) = s^2 + (9/20)s + 1/20
F First terms s^2
O Outside terms (1/5)s OR (4/20)s
I Inside terms (1/4)s OR (5/20)s
L Last terms 1/20